OpenStudy (anonymous):

(3s/2t)-2 -How do you simplify this?

OpenStudy (anonymous):

I dont think you can simplify that much. But you can write it as: \[0.5(3s*y-4)\]

OpenStudy (anonymous):

I dont think that is right. I got this problem from a law of exponent sheet

OpenStudy (welshfella):

is it (3s/2t)^(-2) ?

OpenStudy (anonymous):

\[\left( \frac{ 3s }{ 2t } \right)^{-2}=\frac{ 1 }{ \left( \frac{ 3s }{ 2t } \right)^2 }=\left( \frac{ 2t }{ 3s } \right)^2=?\]

OpenStudy (welshfella):

-2 being the exponent

OpenStudy (anonymous):

 thats how it looks